OpenStudy (anonymous):

Robby skateboards 0.50 blocks to his friend's house in 1.2 minutes. What is his speed?

OpenStudy (anonymous):

average speed is .5 blocks/1.2 minutes, or .42 blocks/minute.

OpenStudy (anonymous):

Speed = Distance/Time Speed = 0.5/1.2 Speed = 5/12 OR 0.4167